OpenStudy (anonymous):

Help with Simplifying Radicals 1. *square root* 49a^8 2. 2 *square root* 10 *times* 3 *square root* 12 3. *square root * 14q *times* 2 *square root* 4q 4. 4/ *square root* 11 i really hope you understand me please help me ❤❤❤❤❤❤

OpenStudy (john_es):

1. \[\sqrt{49 a^8}=\sqrt{7^2(a^4)^2}=7a^4\] 2.\[\sqrt{10}\sqrt[3]{12}=10^{1/2}12^{1/3}=10^{3/6}12^{2/6}=\sqrt[6]{10^312^2}\] 3.\[\sqrt{14q}\sqrt{4q}=\sqrt{64q^2}=8q\] 4.\[4\sqrt{11}=\sqrt{11\cdot4^2}\] May be I missunderstand some of your questions. Tell me if that is the case.

OpenStudy (anonymous):

well actually only one of them is right lol let me see if i can do it with the butttons on here

OpenStudy (anonymous):

thats number 3. 2\[2 \sqrt{10}\times 2 \sqrt{4q}\]

OpenStudy (anonymous):

I mean number 3

OpenStudy (john_es):

3) \[2\sqrt{10}2\sqrt{4q}=4\sqrt{40q}=4\sqrt{2^2\cdot10}=8\sqrt{10}\] I suppose it is right now. Check it :).

OpenStudy (anonymous):

4. \[\sqrt{14q}\times 2 \sqrt{4q}\]

OpenStudy (anonymous):

and for number 3 that answer is not in my options .-.

OpenStudy (john_es):

52 John_ES 0 Umm, the factor 2 is the exponent of the radix or a constant factor? If it is the exponent, then, \[\sqrt{10}\sqrt{4q}=\sqrt{40q}=2\sqrt{10q}\]

OpenStudy (anonymous):

im so sorry i suck at this :c the 2 for what number ? lol im sorry

OpenStudy (john_es):

I say the number 2 in front of the square root \[2\sqrt{}\] As you write is a number that multiplies the square root. But in this other form \[\sqrt[2]{}\] is the exponent. Which of these forms are in your exercise?

OpenStudy (john_es):

May be it should be better if you can capture an image of the exercise.

OpenStudy (anonymous):

oh no thats a regular number just like the top one

OpenStudy (anonymous):

i don't have a screenshot on my laptop

OpenStudy (john_es):

Sorry for the absence. For the fourth question, \[\sqrt{14q}\times2\times\sqrt{4q}=4q\sqrt{14}\]. (You can take a screenshot with the button Imp. Pant, just in the upper-right part of the keyboard. Then, you can use paint to paste the content of the notepad and save it as png)